Let f(x) = |2x − 10| + 2 .
a) Show that f is continuous at x = 5
b) Show that f is not differentiable at x = 5
1
Expert's answer
2020-12-08T06:48:35-0500

a)

Let "\epsilon>0" be given. We must find a "\delta=\delta(\epsilon) >0" such that if"|x-5|<\delta" then, "|f(x)-f(5)|<\epsilon."

Now,


"|f(x)-f(5)|=||2x-10|+2-2|=||2x-10||=2|x-5|<2\delta =\epsilon \, \text{if} \,\delta=\epsilon/2\\
\implies 
|f(x)-f(5)|<\epsilon"

Hence, given "\epsilon>0," choose "\delta=\epsilon/2." Then, "|f(x)-f(5)|<\epsilon" whenever "|x-5|<\delta," so that f is continuous at "x=5."


b)


To show that f is not differentiable at x=5, we must show that the left hand derivative is not equal to the right hand derivative at 5. That is, "f'(5^-) \neq f'(5^+)."

"f'(5^-)=\lim_{x \to 5^-}\frac{f(x)-f(5)}{x-5}=\lim_{h \to 0}\frac{f(5-h)-2}{5-h-5}=\lim_{h \to 0}\frac{2+2h-2}{-h}=\lim_{h \to 0}\frac{2h}{-h}=-2"

"f'(5^+)=\lim_{x \to 5^+}\frac{f(x)-f(5)}{x-5}=\lim_{h \to 0}\frac{f(5+h)-2}{5+h-5}=\lim_{h \to 0}\frac{2+2h-2}{h}=\lim_{h \to 0}\frac{2h}{h}=2"


"-2\neq2 \implies f'(5^-)\neq f(5^+)"

Hence, f is not differentiable at x=5.